OpenStudy (anonymous):

Find the integral of sin^2xcos^4xdx

OpenStudy (anonymous):

\[\int sin^2xcos^4x dx\] Now make them linear I mean \[\int \frac{4(sin^2x.cos^2x) cos^2x}{4}\]

OpenStudy (anonymous):

\[\frac{1}{4}\int sin^22x \times \frac{(cos2x - 1 )}{2}\]

OpenStudy (anonymous):

\[\frac{1}{8}\int sin^22x cos2x - sin^22x dx \]

OpenStudy (sriram):

cos^2 x=[cos2x+1]/2 not [cos2x-1]/2

OpenStudy (anonymous):

Then first integral can be solved by t = sin2x substitution

OpenStudy (anonymous):

\[\frac{1}{8} ( \frac{1}{2} \times \frac{1}{3} t^3 ) + \frac{1}{8}\int sin^22xdx\]

OpenStudy (anonymous):

Now you can convert sin^2x into linear too

OpenStudy (anonymous):

\[sin^22x = \frac{1 - cos4x}{2}\]

OpenStudy (anonymous):

\[\frac{1}{8 * 2 *3}sin^32x + \frac{1}{8*2}x - \frac{1}{8*2*4}sin4x +C\]